### v2

# DCAT: Dual Cross-Attention Fusion for Disease Classification in Radiological Images with Uncertainty Estimation

Abstract
Accurate and reliable image classification is crucial in radiology, where diagnostic decisions significantly impact patient outcomes. Conventional deep learning models tend to produce overconfident predictions despite underlying uncertainties, potentially leading to misdiagnoses. Attention mechanisms have emerged as powerful tools in deep learning, enabling models to focus on relevant parts of the input data. Combined with feature fusion, they can be effective in addressing uncertainty challenges. Cross-attention has become increasingly important in medical image analysis for capturing dependencies across features and modalities. This paper proposes a novel dual cross-attention fusion model for medical image analysis by addressing key challenges in feature integration and interpretability. Our approach introduces a bidirectional cross-attention mechanism with refined channel and spatial attention that dynamically fuses feature maps from EfficientNetB4 and ResNet34 leveraging multi-network contextual dependencies. The refined features through channel and spatial attention highlights discriminative patterns crucial for accurate classification. The proposed model achieved AUC of 99.75%, 100%, 99.93% and 98.69% and AUPR of 99.81%, 100%, 99.97%, and 96.36% on Covid-19, Tuberculosis, Pneumonia Chest X-ray images and Retinal OCT images respectively. The entropy values and several high uncertain samples give an interpretable visualization from the model enhancing transparency. By combining multi-scale feature extraction, bidirectional attention and uncertainty estimation, our proposed model strongly impacts medical image analysis.
Keywords deep learning  $\cdot$ feature fusion  $\cdot$ cross-attention  $\cdot$ classification  $\cdot$ uncertainty quantification
1 Introduction
Deep learning (DL) has gained significant traction in medical imaging due to its ability to automate feature extraction, enhance training efficiency, and deliver superior performance across tasks such as classification  (Affonso et al., 2017; Sahu and Kashyap, 2025; Mohanty et al., 2024; Thiranjaya et al., 2024), segmentation  (Wang et al., 2021; Han et al., 2024), image generation  (Pezoulas et al., 2024; Lang et al., 2024), object detection  (Ragab et al., 2024; Morita et al., 2024) etc. In particular, automated disease classification in radiology has emerged as a promising clinical diagnosis and treatment planning tool. However, accurate classification remains critical, as errors can lead to misdiagnosis, delayed treatment, or unnecessary interventions. Despite its successes, DL-based lung disorder detection faces notable challenges, including false positives, false negatives, and the complexity of radiological images, which often contain overlapping abnormalities, subtle disease patterns, and intricate anatomical structures  (Wang et al., 2017; Rajpurkar et al., 2017). Standard Convolutional Neural Networks (CNNs) often fall short in addressing these challenges due to their limited ability to capture global context, dynamically prioritize multi-scale features, and integrate contextual relationships crucial for accurate diagnosis  (Wang et al., 2020; Lin et al., 2017; Ramamurthy et al., 2024). Attention mechanisms have emerged as a powerful tool in DL, enabling models to dynamically prioritize the most informative features for a given input, enhancing adaptability to variations in data distributions. By focusing on less affected features by distribution shifts, attention mechanisms improve model robustness and generalizability across diverse domains  (Oktay et al., 2018; Liu et al., 2024).
Additionally, attention weights can offer interpretable insights into the model’s confidence and decision-making process, contributing to more calibrated predictions. In medical image analysis, where data variability and complexity are significant challenges, attention-based feature fusion proves particularly advantageous  (Hou et al., 2021; Zheng et al., 2023). Recent studies,  (Valanarasu et al., 2021; Li et al., 2022; Jiang et al., 2021), have explored transformer-based approaches, yet self-attention mechanisms often struggle to capture fine-grained local features, leading to difficulties in distinguishing objects of interest from complex backgrounds. Cross-attention feature fusion addresses this limitation by explicitly modeling relationships between different networks or modalities or views, enabling the model to leverage contextual information from one source to enhance the processing of another  (Ates et al., 2023; Song et al., 2022; Zheng et al., 2023). Unlike traditional fusion methods like concatenation or element-wise addition, cross-attention captures complex dependencies and dynamically weights contributions from each modality, resulting in more informative and nuanced representations. This capability has been demonstrated to improve performance in multi-model and multi-view task significantly  (Wang et al., 2023a; Liu et al., 2024), making cross-attention a promising technique for advancing medical image analysis and other complex visual recognition challenges.
Interpretability and uncertainty quantification (UQ) are pivotal challenges  (Tjoa and Guan, 2020; Begoli et al., 2019) in the application of DL models to radiology image classification  (Faghani et al., 2023; Huang et al., 2024). The inherent complexity of these models often limits transparency, making it difficult for clinicians to understand how decisions are made. This is particularly problematic in radiology, where high inter-patient variability, differences in imaging protocols, and noise can lead to inconsistent classification outcomes. Moreover, traditional classification networks frequently fail to provide uncertainty estimates, resulting in overconfident misclassifications undermining their reliability in clinical practice. So, quantifying uncertainty is essential for assessing the confidence of model predictions, enabling radiologists to weigh the risks and benefits of diagnostic decisions.
In medical image analysis, recent advancements, such as Medical Transformer  (Valanarasu et al., 2021), GasHis Transformer  (Chen et al., 2022), and TransFuse  (Zhang et al., 2021) have demonstrated the efficacy of combining CNNs with Transformers to capture both global and local features, thereby enhancing performance in tasks like disease detection and segmentation. These approaches emphasize the simultaneous learning of local spatial structures and long-range dependencies to reduce noise and improve feature representation. In contrast, our work introduces a novel Dual Cross-Attention (DCAT) fusion mechanism with enhanced Convolution Block Attention Module (CBAM). By integrating cross-attention between EfficientNetB4 and ResNet34 architectures, our model effectively captures multi-scale features, preserving low-level spatial textures and high-level semantic representations. This approach leverages attention-guided multi-scale global-local fusion to enhance feature representation and address challenges in medical image classification. We incorporate MC Dropout during inference to improve robustness and quantify prediction uncertainty. Validated on four medical image datasets, our proposed DCAT fusion model demonstrates superior performance in lung disease classification from chest X-ray images, and retinal disease classification from OCT images.
1.1 Contribution
The work proposes a new, simple, effective DL model considering its uncertainty for medical image classification.
  1. 1.
    We introduce a cross-attention mechanism that dynamically fuses feature maps from two efficient networks, EfficientNetB4 and ResNet34. DCAT leverages the strength of both the architectures, a bidirectional feature exchange significantly improving feature representation for accurate medical image classification.
  2. 2.
    The integration of channel and spatial attention mechanism refines feature selection. Combining hierarchical cross-attention fusion with enhanced CBAM, captures fine-grained details and high-level semantic information.
  3. 3.
    Incorporating uncertainty quantification provides confidence estimates for each prediction demonstrating the reliability with uncertainty measures for enhancing trustworthiness in clinical decision-making.
The remaining paper follows the given hierarchy: the related works, conventional CNN-based and transformer-based approaches, Attention Mechanism & Feature Fusion, and employed uncertainty quantification are discussed in Section 2, while the proposed model details are provided in Section 3. The details on experiments are provided in Section 4 followed by experimental result and analysis in Section 5 along with the in-depth discussion of attained results are presented in Section 6. Finally, the conclusion and future work are mentioned in Section 7.
Figure 1: (a) Random example images from the four datasets. From top Tuberculosis chest X-ray, Covid Chest X-ray, Pneumonia Chest X-ray, Retinal OCT images, (b) Visualization of class distribution for each dataset.
2 Related work
Deep Learning in Radiology.
A wide variety of DL models have shown promising outcomes in complicated diagnostic areas across radiology  (Borah et al., 2024a; Fritz et al., 2023; Borah et al., 2023), pathology  (Ahmed et al., 2022; Unger and Kather, 2024), ophthalmology Jin and Ye (2022); Kazi (2024); Borah et al. (2024b), dermatology, and so on. CNN has been the cornerstone of medical image analysis due to their ability to capture spatial hierarchies in data. Transformer models, introduced by Ashish (2017), have revolutionized the field by effectively handling long-range dependencies through self-attention mechanisms. Models like TransUNet  (Chen et al., 2024) and EG-UNETR  (Cheng et al., 2024) combine Transformers with CNNs to capture global semantic information while preserving local details. However, Transformer-based models often struggle with computational efficiency and memory constraints, particularly when processing high-resolution medical images and when there is a shortage of data. Additionally, they may fail to preserve fine-grained details, limiting their effectiveness in tasks requiring precise localization. Hybrid models that integrate CNNs and Transformers have emerged as a promising solution, leveraging the strengths of both architectures. However, the computational inefficiency of Transformer-based models and their limited ability to preserve fine-grained details hinder their practical application.
Attention Mechanism & Feature Fusion.
Attention mechanisms have emerged as a powerful tool in computer vision. Attention mechanisms can be broadly categorized into channel attention, spatial attention, and cross attention, each addressing specific aspects of feature representation. Channel attention mechanisms, such as Squeeze-and-Excitation (SE) networks  (Hu et al., 2018) and selective kernel attention  (Chen et al., 2017) allow models to recalibrate channel-wise feature responses, emphasizing important features adaptively. Spatial attention mechanisms, on the other hand, focus on identifying salient spatial regions within feature maps  (Woo et al., 2018; Dai et al., 2017). Techniques like CBAM  (Woo et al., 2018) and deformable convolution  (Dai et al., 2017) have significantly improved tasks requiring precise localization, such as object detection and segmentation. Self-attention mechanisms, popularized by Transformers  (Zhao et al., 2020), capture global dependencies by computing attention weights across all spatial positions  (Wang et al., 2023b). Variants like Swin Transformers  (Liu et al., 2021), Focal Transformers  (Yang et al., 2021), and HRFormer  (Yuan et al., 2021) have extended this capability to process multi-scale information effectively. However, these approaches often neglect to adjust channel dimensions, restricting their capacity to fully leverage channel-wise feature interactions. Cross-attention mechanisms have gained traction for their ability to model relationships between modalities or feature maps. Unlike self-attention, which operates within a single feature space, cross-attention enables exchange between distinct feature representations, enhancing the model’s ability to capture contextual dependencies. For instance,  (Schlemper et al., 2019) introduced an attention gate model for medical image analysis, which suppresses irrelevant regions while highlighting task-specific features. Similarly,  (Guan et al., 2020) proposed a three-branch attention-guided CNN for thorax disease identification, integrating global and local cues to improve performance. These studies underscore the potential of attention mechanisms in addressing challenges such as noise suppression and feature refinement. Despite advancements in attention mechanisms, several limitations persist, particularly in cross-attention multi-network feature fusion. While channel and spatial attention mechanisms have been studied extensively, their integration with cross-attention remains insufficiently explored. Finally, the lack of uncertainty quantification in attention limits their reliability in critical applications like medical imaging.
Uncertainty Quantification.
UQ techniques used in DL have gained significant attention over the last couple of years in medical data analysis  (Faghani et al., 2023; Fayjie et al., 2024; Bai et al., 2024). By identifying uncertain samples, models can flag complex cases for further investigation, enhancing reliability instead of overconfidently making wrong predictions. A review of the recent advancements in uncertainty quantification was presented in  (Zou et al., 2023) which covered different approaches and discussed various applications. Moreover, in  (Gawlikowski et al., 2023) factors that cause uncertainty are highlighted, including variability in real-world situations, errors in measurement systems, model structure, training data, and unknown data. Uncertainty in DL can be categorized as aleatoric and epistemic uncertainty  (Kendall and Gal, 2017). Aleatoric uncertainty arises from inherent noise within the data while epistemic uncertainty stems from model limitations. Recent studies have explored various UQ techniques to improve model reliability in medical applications. For instance, Monte Carlo (MC) Dropout has been widely adopted for its simplicity and effectiveness. Nair et al. demonstrated its utility in multiple sclerosis detection and segmentation showing improved performance and uncertainty estimation  (Nair et al., 2020). Similarly, in Chagas et al. (2023) applied MC-Dropout and test-time augmentation for classifying membraneous nephropathy showing the importance of correlation between prediction variance and accuracy. While UQ has been extensively studied in traditional DL models, its study with attention mechanism lacks explicit uncertainty modeling and remains underexplored. While ensemble learning methods  (Dutschmann et al., 2023; Fayjie et al., 2024) have advanced UQ, they are computationally expensive in resource-constrained environments. This gap motivates our work, to choose and integrate MC-Dropout for UQ into our methodology and evaluate DCAT fusion model for enhanced predictive confidence while maintaining reliability.
Figure 2: (a) t-SNE plots for visualization and comparison of distributions across datasets to identify similarities (overlap) and differences between (a) Covid-19 original dataset with four classes vs. Pneumonia Chest X-ray dataset, (b) TB vs. Pneumonia Chest X-ray dataset
3 Materials and Methodology
3.1 Materials
In our image classification task, we evaluated the proposed DCAT model using four publicly available medical imaging datasets, Covid-19 chest X-ray  (Chowdhury et al., 2020), Tuberculosis chest X-ray  (Rahman et al., 2020), Pneumonia Chest X-ray  (Kermany, 2018), OCT  (Kermany, 2018) images. Figure 1(a) and (b) show some randomly selected samples from four datasets and their class-wise distribution respectively. We randomly split the data into two main sets, where 80% of the data was used for training and the remaining 20% was used as test sets. Two important points to be considered regarding datasets in this work. In the Covid-19 chest X-ray dataset, the original datasets contain four classes (Covid-19, Lung-Opacity, Viral Pneumonia, and Normal). However, Viral Pneumonia, and Normal classes have similar samples as the Pneumonia chest X-ray datasets. To prevent data leakage and overlapping, we discard these two classes and have considered only the Covid-19 and the Lung Opacity classes from the dataset. The t-SNE plot in Figure 2 (a) and (b) gives the visualization and comparison of distributions across datasets to identify similarities (overlap) and differences.
Figure 3: Overview of the proposed DCAT fusion mechanism. The proposed mechanism starts with acquiring chest X-ray and OCT images. The proposed model comprised of local and global feature learning by EfficientNetB4 and ResNet34, cross-attention fusion, and predictive uncertainty estimation. Finally, we have uncertainty estimation quantifying the uncertainty of the model.
3.2 Proposed Methodology
Overview of the Framework
This paper proposes a novel cross-attention fusion called the DCAT model that integrates multi-scale feature extraction using EfficientNetB4 and ResNet34 backbones. As illustrated in Figure 3, the proposed architecture comprises four components: dual-branch feature extraction with EfficientNetB4 and ResNet34, hierarchical multi-scale and cross-attention fusion mechanism, refined features using enhanced CBAM attention mechanism, and classification and predictive uncertainty estimation.
EfficientNetB4, part of the Efficient family introduced by Tan and Le (2019) in 2019 employs a compound scaling method that uniformly scales depth, width, and resolution to optimize performance with fewer parameters. This scaling enables the network to capture fine-grained details and high-level semantic information. ResNet34, part of the Residual Network (ResNet) family introduced by He et al. (2016), addresses the vanishing gradient problem through its residual learning framework, enabling the training of deep networks. The architecture extracts hierarchical features, with early layers capturing low-level details and deeper layers capturing high-level semantics. Its residual connections make it robust to over-fitting, even with limited training data. Combining these architectures, the proposed model leverages their complementary strengths in feature representations.
3.3 Feature Extraction
We consider a medical image classification problem where each sample consists of an input image ($x$) and an associated label $y\in{1,2,....,N_{cl}}$, with $N_{cl}$ representing the total number of classes. The feature extraction process is defined using a feature extractor $F_{ex}$, where $ex\in{1,2,....,N}$ denotes the two pretrained feature extractors used in the study. For each pretrained feature extractor, the extracted features are defined as,
$X^{ex}_{img}=F_{ex}(x)$
,where $X^{eff}_{img}$ and $X^{r}_{img}$ represent the feature maps extracted from EfficientNetB4 and ResNet34, respectively. These feature maps capture distinct characteristics of the input image at multiple scales, providing a rich representation for subsequent processing.
3.4 Hierarchical Multi-Scale Feature Extraction
To improve the representation ability of the extracted features, many researchers designed multiple networks to capture more features  (Li et al., 2021; Xu et al., 2024). Huo et al. (2024) constructed a three-branch hierarchical multi-scale feature fusion network that integrates global and local features. In our work, to improve the representation ability of the extracted features, we employ a hierarchical multi-scale feature extraction strategy using EfficientNetB4 and ResNet34. Feature maps extracted from specific layers of these networks are combined to form a multi-scale representation.
The EfficientNetB4 network processes the input image $x$ through a series of Convolutional blocks. Feature maps extracted from EfficientNetB4 Block 3 and 4 have spatial sizes of $28\times 28$ and $14\times 14$, respectively. Similarly, the ResNet34 network processes the input through a series of residual layers, with feature maps extracted from ResNet Layer 2 and 3 having spatial sizes of $28\times 28$ and $14\times 14$, respectively. These feature maps are denoted as,
[…]
The multi-scale feature extraction process produces two sets of concatenated feature maps,
$F^{1}=[X^{eff}_{1},X^{r}_{1}]\in\mathbb{R}^{2C_{1}\times 28\times 28},F^{2}=[X^{eff}_{2},X^{r}_{2}]\in\mathbb{R}^{2C_{1}\times 14\times 14}$
where $F^{1}$ represents fine-grained features at a spatial resolution of $28\times 28$, and $F^{2}$ represents high-level features at a spatial resolution of $14\times 14$. These multi-scale features are then passed to the cross-attention fusion module for further processing.
3.5 Cross Attention mechanism
The attention mechanism computes attention weights for each feature map, identifying its relevance to the classification task. Given an image feature map $F\in\mathbb{R}^{C\times H\times W}$, where $C,H,$ and $W$ denote the number of channels, height, and width of the feature tensor, respectively, our approach employs cross-attention fusion followed by channel and spatial attention modules. Unlike conventional methods, our framework first facilitates feature exchange between EfficientNetB4 and ResNet34 architectures, Figure 3, ensuring that the attention mechanism is guided by multi-network contextual dependencies rather than isolated feature representations. Building on prior works such as  (He et al., 2020; Li et al., 2019; Xie et al., 2021), which proposed channel and spatial attention mechanisms to explore discriminative features, our module ensures the network prioritizes disease-specific features by leveraging cross-attention between architectures, enabling context-aware attention refinement. Unlike traditional CBAM  (Woo et al., 2018) and SENet  (Hu et al., 2018), which apply sequential channel and spatial attention within a single network, our enhanced mechanism integrates cross-network feature interactions. This approach enhances spatial understanding and multi-scale feature integration, addressing the limitations of single-network attention mechanisms.
Let $X^{eff}_{img}\in\mathbb{R}^{C\times H\times W}$ represent the feature map from EfficientNetB4, and $X^{r}_{img}\in\mathbb{R}^{C\times H\times W}$ represent the feature map from ResNet34. To enable bi-directional attention fusion, we compute $\text{Query}(Q),\text{Key}(K)$, and $\text{Value}(V)$ representations of both networks. For EfficientNetB4, Q termed as $Q_{eff}$ is derived from ResNet34 features and are the transformed EfficientNetB4 features (shown in Figure 3). K and Q termed as $K_{eff}$ and $V_{eff}$ respectively are computed from EfficientNetB4 features.
[…]
$Q_{eff}=W_{q}X^{r}_{img},K_{eff}=W_{k}X^{eff}_{img},V_{eff}=W_{v}X^{eff}_{img}$
For ResNet34, Q termed as $Q_{r}$ is derived from EfficientNetB4 features and are the transformed ResNet34 features (shown in Figure 3). K and Q, termed as $K_{r}$ and $V_{r}$ respectively, are computed from the ResNet34 features.
[…]
$Q_{r}=W_{q}X^{eff}_{img},K_{r}=W_{k}X^{r}_{img},V_{r}=W_{v}X^{r}_{img}$
Here, $W_{q}$, $W_{k}$, and $W_{v}$ are learnable weight matrices that project input features into attention space. The attention weights are computed using the scaled dot-product attention mechanism. For EfficientNetB4 attending to ResNet34, the attention is calculated as,
$\text{Attention}(Q_{eff},K_{r},V_{r})=\text{Softmax}\left(\frac{Q_{eff}K_{r}^{T}}{\sqrt{d_{k}}}\right)V_{r}$ (1)
where $Q_{eff}K_{r}^{T}$ measures the similarity between the queries and keys, and $\sqrt{d_{k}}$ is a scaling factor that stabilizes gradients. Similarly, for ResNet34 attending to EfficientNetB4, the attention is computed as,
$\text{Attention}(Q_{r},K_{eff},V_{eff})=\text{Softmax}\left(\frac{Q_{r}K_{eff}^{T}}{\sqrt{d_{k}}}\right)V_{eff}$ (2)
where $Q_{r}K_{eff}^{T}$ measures the similarity between the queries and keys, and $\sqrt{d_{k}}$ is a scaling factor that stabilizes gradients. The final fused feature map $F_{fusion}$ is obtained by combining the attention outputs from both the directions,
$F_{fusion}=\text{Attention}(Q_{eff},K_{r},V_{r})+\text{Attention}(K_{eff},Q_{r},V_{eff})$ (3)
This bidirectional fusion ensures that the final feature representation incorporates complementary information from both networks, enhancing the model’s ability to capture subtle patterns and relationships in the data.
(a) Channel Attention Module
(b) Spatial Attention Module
Figure 4: (a) Channel-attention and (b) spatial attention module as part of the CBAM used in this study. The proposed DCAT fuses feature maps from the two pretrained models before applying channel and spatial attention, enabling our proposed model to leverage information from the two networks and enhancing the feature representations for the classification task.
3.6 Feature Refinement and Classification
Following the cross-attention fusion, we refine the feature maps using an enhanced version of the CBAM attention mechanism (Figure 4). By leveraging channel and spatial attention mechanisms, our refined CBAM enables the model to attend to the channel and spatial axes, respectively.
3.6.1 Channel Attention Mechanism.
The channel attention mechanism computes attention weights for each channel in the fused feature map $F_{fusion}\in\mathbb{R}^{C\times H\times W}$ where $C,H,$ and $W$ represent the number of channels, height, and width, respectively. To aggregate spatial information, both average pooling and max pooling are applied to $F_{fusion}$,
$\text{AvgPool}(F_{\text{fusion}})=\frac{1}{H\times W}\sum_{i=1}^{H}\sum_{j=1}^{W}F_{\text{fusion}}(i,j)]$ (4)
$\text{MaxPool}(F_{\text{fusion}})=\max_{i\in H,j\in W}F_{\text{fusion}}(i,j)$ (5)
The pooled features are combined and passed through a shared multi-layer perceptron (MLP) with weights $W_{1}$ and $W_{2}$, followed by ReLU activation and a sigmoid function ($\sigma$),
$M_{c}=\sigma(W_{2}ReLU(W_{1}(\text{AvgPool}(F_{fusion})+\text{MaxPool}(F_{fusion})))$ (6)
Here, $M_{c}\in\mathbb{R}^{C\times 1\times 1}$ represents the channel attention weights, which highlight the most important channels. The refined feature map $F_{refined}$ is obtained by multiplying $M_{c}$ with $F_{fusion}$.
3.6.2 Spatial Attention Mechanism.
The spatial attention mechanism computes attention weights for each spatial location in the feature maps. Both average pooling and max pooling are applied along the channel dimension, and the results are concatenated. A convolutional layer with weights $W_{s}$ is applied, followed by a sigmoid function,
$M_{s}=\sigma(W_{s}\times[\text{AvgPool}(F_{fusion}),\text{MaxPool}(F_{fusion})])$ (7)
Here, $M_{s}\in\mathbb{R}^{1\times H\times W}$ represents the spatial attention weights. The final refined feature map is obtained by applying both channel and spatial attention mechanisms. This joint refinement process ensures that important channels and spatial regions are emphasized, enhancing the discriminative power of the feature. This approach significantly improves feature representation, making it particularly effective for medical image analysis tasks.
3.6.3 Feature pooling and classification.
The refined feature maps $F_{refined}\in\mathbb{R}^{C\times H\times W}$ are passed through an adaptive average pooling layer to reduce their spatial dimensions to a fixed size, ensuring compatibility with the fully connected layer regardless of the input image size,
$F_{pooled}=\text{AdaptiveAvgPool}(F_{refined})$ (8)
where $F_{refined}\in\mathbb{R}^{C\times 1\times 1}$ represents the pooled feature map. The pooled features are then flattened into a vector and passed through a fully connected layer with weights $W_{c}$ and bias $b$,
$y=\text{Softmax}(W_{c}F_{pooled}+b)$ (9)
Here, $y$ represents the output probability distribution over the possible class labels. The sigmoid activation function ensures that the output values sum to one, providing a probabilistic interpretation of the model’s predictions. The class with the highest probability is selected as the predicted label $\hat{y}$.
3.7 Model Uncertainty Calculation.
MC-Dropout.
A widely used technique for quantifying uncertainty and introduced by Gal and Ghahramani  (Gal and Ghahramani, 2016), this method treats dropout regularization as a Bayesian approximation, enabling scalable learning of predictive distributions. Traditionally, dropout is applied during training to prevent over-fitting by randomly deactivating neurons, which reduces reliance on specific training data patterns. MC Dropout extends this concept by activating dropout during training and testing. This approach involves running multiple stochastic forward passes on the same input, effectively generating a predictive distribution. The resulting softmax outputs are averaged to form the final prediction while also reflecting the model’s uncertainty.
Consider a trained network model, $f_{\theta}(\cdot)$ where $f$ is the network and $\theta$ is the learnable parameters with an active dropout. Suppose a new test $x,y$ is to be predicted, where $x$ is the array of images and $y$ is the corresponding label. It will be passed through the network with active dropout $M$ times (i.e. no. of forward passes). Thus, the predictive posterior distribution can be approximated to generate the final prediction as follows:
$p(y|x)=\frac{1}{M}\sum_{m=1}^{M}{p}_{m}$ (10)
where ${p}_{m}$ denotes the output of the softmax function at pass $m\in$ $1,...,M$.
${p}_{m}=\text{Softmax}(f_{{\theta}_{m}}(x))$ (11)
Following the work of  (Gal and Ghahramani, 2016), we enable the dropouts during inference. Our experiment computes the posterior probability distribution over the trained network weights given the input sample and the corresponding ground truth. 100 forward passes are performed to quantify uncertainty, and an average is taken to produce the final prediction. We used a dropout with a rate of 0.5 during training.
4 Experiments
4.1 Implementation details
We choose Pneumonia, Covid-19, Tuberculosis, and Normal for chest X-ray images belonging to three different datasets and CNV, DME, DRUSEN, and Normal for retinal OCT images as the categories of classification task. Our chest X-ray images are collected from public repositories. We implement the proposed mechanism and perform experiments using the PyTorch library. Training and evaluations have been accomplished using NVIDIA GeForce RTX 3060 Ti inbuilt GPU. The batch size for training is set to 32. The maximum epochs is set to 200 and the learning rate is initialized to $43e-$ 4 with a weight decay of $41e-$ 4. We choose Adam as the optimizer for the training stage. The whole training process takes 2h30min and the testing time for a per radiology image is approximately 0.015s.
4.1.1 Evaluation metrics
For a comprehensive overview and evaluation of the performance of the proposed model, we employed standard performance metrics like accuracy, AUC is the area under the ROC curve (AUROC), Area Under the Precision-Recall (AUPR), precision, recall, specificity, F1-score, Matthews Correlation Coefficient (MCC) and Cohen’s Kappa. As our dataset is class-imbalanced, we use the AUPR as a metric that evaluates classification performance. AUC, AUPR focus on positive class performance. MCC and Cohen’s Kappa are both metrics used to evaluate classification performance, but they have distinct differences in handling agreement and correlation. Since our datasets are class imbalanced, we considered MCC for evaluating the classifier performance. The Kappa coefficient, is used to measure the classification accuracy of the model and can also be used for consistency testing.
$\text{Accuracy}=\frac{TP+TN}{TP+FP+TN+FN}$ (12)
$\text{Precision}=\frac{TP}{TP+FP}$ (13)
$\text{Recall}=\frac{TP}{TP+FN}$ (14)
$\text{Specificity}=\frac{TN}{TN+FP}$ (15)
[…]
$\text{MCC}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+TN)(TP+FN)(TN+FP)(TN+FN)}}$ (17)
$\text{Kappa}=\frac{p_{0}-p_{e}}{1-p_{e}}$ (18)
Here, $TN,TP,FN,FP$ are true negative, true positive, false negative and false positive respectively. In Cohen Kappa, $p_{0}$, and $p_{e}$ represent observed agreement, and expected agreement by chance.
4.1.2 Uncertainty evaluation metrics
Entropy.
Entropy is directly related to the confidence of a model’s prediction. When entropy is high, it signals that the model is uncertain and that it might benefit from additional information and vice versa. It’s a measure of variability associated with random variables  (Shannon, 1948). Entropy facilitates transparency by quantifying uncertainty, making it easier for clinicians to make informed decisions. Since medical imaging deals with complex, noisy, and ambiguous data, images like X-rays can have varying levels of quality and precision. In the context of predictive uncertainty, entropy measures the uncertainty of a probability distribution over possible outcomes and is defined mathematically as -
$H(p(y|x))=-\sum_{c=1}^{C}p(y|x)logp(y|x),$ (19)
Here, C is the number of classes.
Table 1: Performance comparison of our proposed DCAT applied to four datasets at evaluation stage with uncertainty quantification. HUS $\rightarrow$ High Uncertainty Sample
Datasets Accuracy Mean Entropy Standard deviation HUS Misclassified Samples
Covid-19 97.41 0.0254 0.1011 3 12
Tuberculosis 99.21 0.0024 0.0353 7 10
Pneumonia 98.95 0.0150 0.0817 5 11
OCT 91.17 0.1603 0.2457 45 36
Table 2: Details of class-wise classification performance of four datasets used in this study. CNV $\rightarrow$ Choroidal neovascularization, DME $\rightarrow$ Diabetic Macular Edema
Datasets Classes Precision Recall Specificity F1-score
Covid-19 Chest X-ray Covid-19 94.63 98.17 98.17 96.37
Lung Opacity 98.86 96.60 96.53 97.72
Pneumonia Chest X-ray Pneumonia 98.29 98.55 98. 63 98.42
Normal 96.14 95.47 95.43 95.80
Tuberculosis Chest X-ray Tuberculosis 99.21 94.74 94.74 96.92
Normal 96.02 99.41 99.41 97.68
Retinal OCT CNV 92.90 90.04 97.39 91.45
DME 96.13 91.24 98.80 93.62
DRUSEN 88.59 83.87 96.98 86.16
NORMAL 87.44 98.43 95.04 92.61
00footnotetext: CNV: Choroidal neovascularization, DME: Diabetic Macular Edema
5 Experimental Results and Analysis
A comprehensive assessment of the proposed model’s performance was conducted and the findings are reported through performance metrics in the subsequent tables. To evaluate the effectiveness of the proposed model from different perspectives, we experimented on four different medical imaging datasets as listed in the earlier section. The experimental results of the four datasets are shown in Table 1, 2 and 3 respectively. The results from our ablation study are shown in Table 4, effectively demonstrating the performance of various individual components within the DCAT on classification performance. We also evaluated the performance with some of the already published works.
As shown in Table 5, demonstrates the effectiveness of our DCAT model. Our proposed model achieved accuracy of 97.41%, 99.21%, 98.95%, and 91.17% for Covid-19 chest X-ray, tuberculosis chest X-ray, pneumonia chest X-ray, and retinal OCT datasets, respectively. The AUROC and PR-curves of the proposed DCAT are shown in Figures 5, for the three chest X-ray datasets, and the retinal OCT dataset respectively.
Table 3: Performance comparison of our proposed DCAT applied to various chest X-ray and retinal OCT images at evaluation stage
Datasets AUC AUPR Precision Recall Specificity F1-Score MCC kappa
Covid-19 99.75 99.81 98.21 97.35 97.50 97.78 94.68 94.67
Tuberculosis 100.0 100.0 100.0 98.68 100.0 99.34 98.83 98.82
Pneumonia 99.93 99.97 99.21 99.34 97.91 99.27 97.35 97.35
OCT 98.69 96.36 91.33 91.17 97.05 91.13 88.27 88.17
Figure 5: AUROC Curve and Precision-Recall Curve. From the top-left: Retinal OCT (a) AUROC Curve (b) Precision-Recall Curve, Covid-19 Chest X-ray (c) AUROC Curve (f) Precision-Recall Curve, Pneumonia Chest X-ray (d)AUROC Curve (g) Precision-Recall Curve, and Tuberculosis Chest X-ray (e) AUROC (h) Precision-Recall Curve
5.1 Ablation Studies
A comprehensive ablation study was conducted to validate and justify the component of the proposed model on the four datasets and the evaluation metrics previously discussed. The analysis supports our design choices, thereby providing valuable insights into each element’s significance in enhancing the performance of the DCAT, and achieving the desired outcomes. Table 4 summarizes the dataset’s results without DCAT mechanism. This shows that each component of the model contributes to the overall performance. Table 5 compares our proposed DCAT with other recently proposed deep learning methods.
Table 4: Performance comparison of EfficientNetB4 models ResNet34 at individual level without cross attention mechanism on the four datasets
ResNet34111The reported results are average performance of DCAT on the four datasets used in the study. HUS: High Uncertainty Sample. EfficientNetB422footnotemark: 2
Datasets Accuracy Entropy Standard deviation Accuracy Entropy Standard deviation
Covid-19 86.99 0.0443 0.0239 88.76 0.0926 0.0172
Tuberculosis 80.94 0.1896 0.0948 85.89 0.2000 0.0968
Pneumonia 81.89 0.0500 0.0467 80.77 0.0756 0.0891
OCT 88.91 0.2646 0.0367 90.18 0.1290 0.2567
Figure 6: Uncertainty distribution visualization of a subset of data from each dataset with all samples and samples with high uncertainty above an uncertainty threshold. From the left: (a) Covid-19 Chest X-ray, (b) Tuberculosis Chest X-ray, (c) Pneumonia Chest X-ray, and (d) Retinal OCT
Table 5: Performance comparison of various deep learning models and our proposed DCAT fusion model applied to OCT, Pneumonia, Tuberculosis, and Covid-19 datasets at the evaluation stage
Methods Datasets Accuracy AUC UQ Mean Entropy Standard deviation HUS
Abdar et al. (2021) Pneumonia 81.57 92.60 Yes  –  –  –
Yang et al. (2023) Pneumonia 94.60 99.10 No  –  –  –
Chhikara et al. (2020) Pneumonia 90.70  – No  –  –  –
DCAT (Ours) Pneumonia 98.95 99.93 Yes 0.0150 0.0817 13
Abdar et al. (2021) OCT 91.40 98.21 Yes  –  –  –
Yang et al. (2023) OCT 77.60 96.20 No.  –  –  –
DCAT (Ours) OCT 91.17 98.60 Yes 0.1603 0.2457 45
00footnotetext: HUS: High Uncertainty Sample
Figure 7: Random high uncertain samples from four datasets with entropy value, predicted probability, and ground-truth. From the left: (a) Covid-19 Chest X-ray, (b) retinal OCT, (c) Pneumonia Chest X-ray, and (d) Tuberculosis Chest X-ray dataset
Figure 8: Visualization of Grad-CAM analysis on chest X-ray and OCT datasets using DCAT from the left: Original chest X-ray and OCT images, their corresponding heat maps, and the Grad-CAM overlays and repeating the same towards the right.
6 Discussion
In the last few decades, the impressive performance of DL methods broadly applied to various fields of application, from engineering to healthcare, is a reason for using such methods in analyzing medical data. However, due to insufficient labeled medical datasets, most DL methods rarely achieve the desired performance. In this article, inspired by previous studies, we therefore propose a cross-attention fusion method called DCAT for classification of various medical image datasets. Unlike most previous fusion models, the proposed DCAT fusion model takes the advantage of a new approach called deep hierarchical dual cross-attention fusion, which can separate different classes within the datasets. This can be visualized by the AUPR and MCC values given in Table 3.
We developed a novel hierarchical DCAT fusion model as stated in Figure 3. As indicated, we hierarchically obtain the features from EfficientNetB4, which is our first feature extractor, and then the second feature extractor, ResNet34. Unlike the previous feature fusion, instead of using original features extracted by pretrained models, we proposed a feature fusion based on the cross attention mechanism. An attention mechanism is utilized to detect the relevant features in each stage and then combine them. In addition, as most previous studies did not consider estimating the uncertainty in their model, we included uncertainty estimation in our experimental study. Unlike prior studies applied with feature fusion or cross attention, including simple dropout or performance evaluation on standard metrics, where we do not get enough information on how uncertain and certain the model is in its predictions and its reliability. We overcome this in our proposed work by showing entropy as a metric of uncertainty evaluation and finding the number of total highly uncertain samples to flag them for additional expert investigation and review. This is a critical component for reliability and interoperability in any DL-based systems in safety-critical applications such as medicine and healthcare. We quantify uncertainties during all the stages of our experimentation. To better reveal the importance and impact of our proposed DCAT, its performance is compared with other methods in Table 5. The visual explanation of the proposed DCAT fusion model using heat map and gradient-weighted class activation mapping (Grad-CAM) techniques is presented in Figure 8. Finally, the probability distribution of the predictions for each class obtained by the proposed DCAT fusion model for four datasets are presented in Table 2, respectively. In Table 1, we show some high uncertainty samples (HUS) and misclassified samples. In Figure 7 we demonstrate high uncertain samples from each dataset based on entropy, ground-truth and predictive probability. These highly uncertain samples can be flagged to the medical experts for further investigation and clarification.
Altogether, the key advantage of using the proposed DCAT is as follows. 1) It considers a hierarchical dual cross attention where each model contributes to the final prediction where it has the advantage of feature fusion by looking closely attending and informing the model what to look at and where to look at, 2) Using uncertainty quantification method to evaluate the prediction of the model and uncertainties during each stages, and number of high uncertain and misclassified samples. 3) Because our datasets are class-imbalanced, the AUC, AUPR and MCC values in Table 3 show that our model is confident in most cases in discriminating between the positive and negative cases while making a few errors.
In medical imaging, translating a diagnosis into clinically actionable information for evaluating the health of a patient, they are guided by their intuitions and knowledge. Here, the reliability will be on the model’s identification and understanding of features that align with medical practice based evidence. Different variations can lead to variations in DL model’s performance, such as specific medical imaging modality, which has its unique challenges, its data availability and tasks, resolution of images, and its preprocessing requirements, information contents, architectural complexity within the model, and clinical variability. So, understanding the requirements and subtle differences in building effective DL solutions. Considerations incorporating such spectrum into the design of model, training, and evaluation would be effective and transferable in clinical applications. Quantifying the uncertainty within model’s predictions offers a more robust and scalable approach to navigating real-world healthcare data complexities. This facilitates in providing a confidence in the interpretation of medical images for clinical decision-making while contributing to the interpretability of the model allowing the clinicians to identify cases that are most uncertain for further investigation and verification. Future work should investigate the implementation of how the uncertainties within the model can be enhanced through building more robust and adaptive models and extending to CT and MRI data.
7 Conclusion
This article proposed a novel, simple, yet effective feature fusion model with uncertainty quantification based on a hierarchical dual cross-attention fusion approach called DCAT considering two pretrained models. The proposed fusion model was validated using three well-known medical chest X-ray and OCT datasets. Our proposed model had successfully captured the spatial relationships between features extracted by two pretrained DL models and obtained high classification performance with the DCAT feature fusion approach. The model generalized the attention mechanism concept as a selective type of feature fusion into the main learning framework. The model yielded high classification performance owing to the hierarchical multi-scale dual feature fusion strategy. We plan to extend our proposed model to the medical image segmentation task and extend the work to enhance the uncertainties within the model for adaptive and robust performance. Future work also includes investigation and extending to pathophysiology, CT, and MRI data.
Declaration of Competing Interest
The authors declare no competing interests or any relationships that could influence the work reported in this paper.